Use the Squeezing Theorem to show that

lim

X approaches to +sum to infinity

Sin x divide by x =0


1
Expert's answer
2021-06-03T03:20:40-0400
"-1\leq\sin x\leq 1"

If "x>0"


"-\dfrac{1}{x}\leq\dfrac{\sin x}{x}\leq \dfrac{1}{x}"

"\lim\limits_{x\to+ \infin}(-\dfrac{1}{x})=0\ and\ \lim\limits_{x\to +\infin}(\dfrac{1}{x})=0"

Therefore by the Squeezing Theorem


"\lim\limits_{x\to + \infin}(\dfrac{\sin x}{x})=0"